Hi can you helo me find the slope & y-intercept?
6x-12y=24?

Answer by nerdybill(7384)

Rearrange:
6x-12y=24
so that it is in "slope-intercept" form of
y = mx + b
where
m is the slope
b is the y-intercept at (0,b)
.
6x-12y=24
-12y=-6x+24
12y=6x-24
y=(6/12)x-(24/12)
y=(1/2)x-2
from inspection then we see that:
slope is 1/2
y-intercept is at (0, -2)